You deposit $2000 in an account earning 2% interest compounded monthly. How much will you have in the account in 10 years?
Answer by mananth(16949)

Principal P = 2000
Amount= A
years=n 10.00
compounded 12 times a year t
Rate = 2.00 0.02
Amount = P*((n+r)/n)^n*t

Amount = = 2000 *( 1 + 0.02 /t)^ 10 * 12
Amount = 2000 *( 1 + 0 )^ 120
2000 *( 1 )^ 120
Amount = 2442.4